# Intrinsic Resolution of Organic Scintillators

### v1

@@ 7 Advanced measurements with the BC-408 plastic scintillator
The 6x6x1 cm BC-408 plastic scintillator plate under test was optically coupled from all sides to four square photomultiplier tubes (R6236-01, Hamamatsu). The monoenergetic electrons beam was produced by an electron spectrometer capable to provide a range of electron energies from 0.7 MeV to 1.7 MeV (0.7, 1.0, 1.3, 1.5, and 1.7 MeV). The energy spread for 1.0 MeV electron beam was measured to be approximately 1.8% FWHM using high-resolution Si(Li) detector. As reported in [7], this value includes contributions from the energy losses of the electron beam as it propagates through air and materials towards the Si detector, as well as the detector’s resolution. The energy spread for protons of 2.8 MeV and 3.4 MeV measured by the Si detector was confirmed to be less than 1% (FWHM).
The measurement scheme involves recording the charges detected by each of the four PMTs in the set-up, denoted $Q_{1},Q_{2},Q_{3}$ and $Q_{4}$. Two combinations of these values are used in the analysis: the sum of all four charges, $Q^{+}=Q_{1}+Q_{2}+Q_{3}+Q_{4}$, and the differential signal $Q^{-}=(Q_{1}+Q_{3})-(Q_{2}+Q_{4})$. The $Q^{-}$ signal provides a measurement of the statistical part, which can be used to extract the internal resolution from the $Q^{+}$ values. In this scheme, the correlated contributions (internal resolution and the contribution of the electron beam energy spread) are mutually cancelled, providing a significant advantage over schemes with a single PMT. Another advantage of the four-PMT scheme is that it reduces the non-uniformity of the effect of single PMT responses by a factor of four. Indeed, it can be shown that in the case the PMTs exhibit non-uniformity $v(p)_{i}$, the resulting contribution will be a quarter of the average, i.e. $v(p)=\frac{\overline{v(p)}}{4}$, where $\overline{v(p)}=\frac{\sum v(p)_{i}}{4}$.
@@
Figure 5: The fit of the statistical energy resolution data from [7] with $\frac{1+v_{1}}{LY\cdot f_{\text{NL}}\cdot E}$ function.
At the next stage of the analysis, we observe that the residual energy resolution values for protons, once the statistical component has been subtracted, are at the level of $10^{-4}$. As the design of the setup is very similar to previous cases, we can expect a similar contribution from the non-uniformity of the response on a single PMT. The contribution from the signal summed from four PMTs will correspondingly be reduced by a factor of 4, providing a contribution at the level of $v(p)\simeq 10^{-4}$. Therefore, we can conclude that the contribution of intrinsic resolution for protons is $\simeq 0$. The fit of the proton energy resolution data is presented in Fig. 6. The best-fit value for $v(p)=(1.00\pm 0.25)\times 10^{-4}$.
Figure 6: The fit of the resolution data for proton beam from [7] with constrained statistical term and free $v(p)$. The smearing of the proton beam is taken into account as a penalty term in the $\chi^{2}$.
@@
Figure 7: The fit of the data for electrons from [7] with constrained statistical term and $v(p)$. The smearing of the electron beam is also taken into account as a penalty term in the fit.
@@ 8 Measurements with the LAB+PPO liquid scintillator
Using our model, we calculated the value of $v_{1}^{\text{int}}$: $v_{\text{int}}=\frac{Q_{1}}{Q}v_{1}^{\text{int}}=\frac{LY\cdot 1}{LY\cdot 0.05\cdot f_{\text{NL}}(E,kB)}v_{1}^{\text{int}}\simeq\frac{20}{0,9}v_{1}^{\text{int}}$, we get $v_{1}^{\text{int}}=(9.0\pm 2.3)\times 10^{-4}$, which corresponds to an intrinsic resolution of $(3.0\pm 0.4)$at 1 MeV 444the values in our paper [1] are different: $v_{1}^{\text{int}}=(11\pm 3)\times^{-4}$ and $R_{\text{int}}(\text{1 MeV})=(3.3\pm 0.4)$%, this is the result of a miscalculation, division by 0.9 instead of multiplication.. The value $f_{\text{NL}}(k_{\text{B}},E_{\text{e}}=0.05\text{ MeV})\simeq 0.9$ is obtained from Fig.3 of [5].
Below, we reanalyse the data using the additional measurement details provided in PhD thesis [6].
@@
The fit of the spectrum obtained using a low-intensity light source is shown in Fig.2.29 of [6]. The data in the plot for two-photoelectrons contribution can be used to evaluate it variance as $v_{2}\equiv\frac{p5}{p4}=(\frac{0.0758}{0.4246})^{2}=0.032$. However, the relative variance of the single photoelectron spectrum is factor 2 larger, i.e. $v_{1}=0.064$, this already contradicts the $v_{1}=0.05$ value used in the analysis.
Other available measurements of the Hamamatsu R6231-100 PMT show the presence of a strong branch of underamplified signals in its single p.e. spectrum (see Fig.7 from [26]). The values of $v_{1}$, obtained for the two PMT samples in the aforementioned paper, are $v_{1}=0.27$ and $v_{1}=0.28$. This feature is invisible in Fig.2.29 of [6] because of the high threshold set in the measurements. We note that, although the article [26] does not present an estimate of measurement errors, accurately determining the parameter $v_{1}$ is a challenging task for spectra with a significant contribution from underamplified signals due to the parameter’s high sensitivity to spectral shape features.
The presence of a strong underamplified signal branch in the single-electron spectrum will also influence the average SER value: the peak value will shift above 1 p.e., while in the analysis in [6] the position of the single p.e. is chosen practically at the peak value ($q_{\text{peak}}=0.217$ while 1 p.e. corresponds to 0.2$\div$0.21 in $V\cdot ns$ scale [6]). The values of $c$ and $v_{1}$ can be reconstructed using (19) and (20).
@@
The fit was performed using the function:
@@
where $f_{\text{NL}}(E)=f_{\text{NL}}(k_{B}=0.0196,E)$, and $\Delta E$ is the energy smearing in the Ge detector. The fit returns:
@@
with $\chi^{2}=32.7/26$ and $\Delta E=(2.2\pm 1.2)$ keV, see Fig. 8. The value of $v(p)$ does not depend on the choice of the set of the parameters $v_{1}$ and $c$, so we can conclude that $v(p)$ is compatible with zero in this experimental setup, and it is $v(p)<6.0\times 10^{-4}$ at 90% C.L. as results from the $\chi^{2}$ profile study. The value of $\Delta E$ is compatible with the stated width of the energy bin on the Ge detector side of 1 keV, the effect of the Ge detector resolution probably contributes to this value as well.
@@
The obtained value $v_{1}^{\text{int}}=(4.8\pm 2.1)\times 10^{-4}$ corresponds to $R_{\text{int}}(\text{1~{}MeV})=(2.2\pm 0.5)$.
@@ 9 The measurement with LAB-based LS using monoenergetic electrons at 1 MeV
In this section, we examine the measurements reported in [9]. In our previous paper [1], we accepted the reported result without critical assessment. However, as the estimated uncertainties seem too low (see Table 3 of the paper [9]), we investigated the reliability of the presented errors. At first glance, it can be seen that the maximum reported uncertainty in the table is $0.11$%, while the resulting uncertainty is $0.06$%. This appears to be the result of a miscalculation, as the absolute uncertainties should be summed quadratically.
@@
The PMT used for the measurements was calibrated using the peak value of the charge spectrum, as the reported resolution of a single- photoelectron (about 20%) is unrealistically good. Although there are no plots of SER in the paper, but a very similar measurement was presented by a subgroup in another paper [21], which provides the opportunity to reanalyse the spectrum. The SER spectrum is presented in Fig.9 of the paper [21]. We observe a cut pedestal and a discrepancy between the mean values of the two fitted Gaussians: $\text{mean}_{1}=123.4$ and $\text{mean}_{2}=264$. If the two Gaussians represent the single electron response and two-electrons responses, then we should have $\text{mean}_{2}=2\cdot\text{mean}_{1}$. It appears that the pedestal position is not being taken into account. A rough estimate can be obtained from $\text{pedestal}=\text{mean}_{1}-(\text{mean}_{2}-\text{mean}_{1})=-17.2$.
@@
In the case of a Poisson distribution of probabilities of detecting a fixed amount of photoelectrons with an average $\mu$, the probability ratio of a two-electron response to a single electron response is given by $\frac{P(2)}{P(1)}=\frac{\mu}{2}$. The ratio of the corresponding areas gives us $\frac{P(2)}{P(1)}=0.348$, and hence we can estimate $\mu\simeq 0.695\pm 0.044$. The uncertainty is estimated from the (unfitted) fraction of signals in the SER valley.
Using the mean and r.m.s. data from the figure, we derive the following: $q_{1}=\frac{(162.5+17.2)\cdot(1-e^{-\mu})}{\mu}=129.3$ and $v(q)=(\frac{96.86}{162.5+17.2})^{2}=0.290$. The calibration parameter $c$ can be calculated using the values from Fig.7: $c=\frac{\text{mean}_{1}}{q_{1}}=0.92$. Similarly, we estimate $v_{1}=0.10\pm 0.01$ using $v(q)=0.290$. The former should be compared to the reported value of $v_{1}=0.04$, which corresponds to a very low 20% SER resolution. Using these values and taking into account the reported LY=3266 p.e., we obtain for statistical contribution of $\delta_{\text{st}}=(1.68\pm 0.08)$(instead of 1.75%), and for $\delta_{\text{s.p.e.}}=(0.55\pm 0.11)$(instead of 0.35%). The uncertainties of these values are dominated by the uncertainty in extracting $\mu$ value.
The recalculated value of the internal contribution at 0.976 MeV is $R_{\text{int}}(0.976~{}\text{MeV})=(1.85\pm 0.18)$%. Note the larger uncertainty and the fact that the central value has hardly changed compared to the values reported in [9] values. Excluding the double count of the non-uniformity contribution from the final result increases the value of the internal contribution to $R_{\text{int}}(0.976)=(2.06\pm 0.14)$%. We will use the latter value to derive $v_{1}^{\text{int}}$ for this measurement. Note that the possible contribution of the internal resolution of the LaBr3 is not subtracted in the estimates of the systematics,
@@ 10 Discussion
All data are consistent with our simple model within the measurements precision. Note, that in a number of cases, we rescaled the reported uncertainties as they appeared to be overestimated. These data do not allow us to draw any conclusions concerning the presence of any unaccounted additional fluctuations.
The values of the parameter $v_{1}^{\text{int}}$ obtained in our analysis, along with results derived in [1] from Borexino publications, are compiled in Table 3.
The data indicate an additional broadening of the energy resolution due to intrinsic resolution, contributing approximately $\simeq(1.5\div 2)$at 1 MeV for all considered scintillators. For the BC408 (EJ200) plastic scintillator, different experiments report IR contributions in the range $\simeq(1.25\div 2.0)$at 1 MeV. The higher values correspond to measurements around 1 MeV, while lower-energy data yield smaller values. This discrepancy may arise from unaccounted systematics in low-energy measurements, particularly since the results in [7] are free of such systematics. Alternatively, it could suggest different underlying processes in the scintillator response below and above 1 MeV, leading to deviations from the phenomenological model above 1 MeV.
@@
The two Borexino measurements for PC-based LS show good agreement. The value reported in the more recent Borexino publication [12] is likely more precise, as its parameterisation incorporates the energy dependence of the IR contribution. The earlier analysis yields an effective value; converting this into a universal parameter would require an accurate determination of the effective N value at the tail of the 14C spectrum. We further note that this result is unaffected by calibration bias and $v_{1}$ uncertainties, since it was obtained using PMT occupancy (number of triggered PMTs per event) rather than total charge. See [1] for detailed measurement methodology.
@@
Finally, experiments with monoenergetic proton sources show no intrinsic smearing effect, a finding that may be significant for detector calibration.
@@
Table 3: The universal parameter $v_{1}^{\text{int}}$ and the corresponding contribution to the energy resolution at 1 MeV, $R_{\text{int}}(\text{1 MeV}$, for organic scintillators considered in this paper. We add also two measurements from Borexino detector, considered in [1].

### v2

@@ 7 Advanced measurements with the BC-408 plastic scintillator
The 6x6x1 cm BC-408 plastic scintillator plate under test was optically coupled from all sides to four square photomultiplier tubes (R6236-01, Hamamatsu). The monoenergetic electrons beam was produced by an electron spectrometer capable to provide a range of electron energies from 0.7 MeV to 1.7 MeV (0.7, 1.0, 1.3, 1.5, and 1.7 MeV). The energy spread for 1.0 MeV electron beam was measured to be approximately 1.8% FWHM using high-resolution Si(Li) detector. As reported in [6], this value includes contributions from the energy losses of the electron beam as it propagates through air and materials towards the Si detector, as well as the detector’s resolution.
The measurements in this study were conducted also with proton beam lines (microbeams) located at the WERC, Japan (see references in [6]). Two specific energies were used, the energy spread for protons of 2.8 MeV and 3.4 MeV measured by the Si detector was confirmed to be less than 1% (FWHM).
The measurement scheme involves recording the charges detected by each of the four PMTs in the set-up, denoted $Q_{1},Q_{2},Q_{3}$ and $Q_{4}$. Two combinations of these values are used in the analysis: the sum of all four charges, $Q^{+}=Q_{1}+Q_{2}+Q_{3}+Q_{4}$, and the differential signal $Q^{-}=(Q_{1}+Q_{3})-(Q_{2}+Q_{4})$. The $Q^{-}$ signal provides a measurement of the statistical part, which can be used to extract the internal resolution from the $Q^{+}$ values. In this scheme, the correlated contributions (internal resolution and the contribution of the electron beam energy spread) are mutually cancelled, providing a significant advantage over schemes with a single PMT. Another advantage of the four-PMT scheme is that it reduces the non-uniformity of the effect of single PMT responses by a factor of four.
Indeed, the variance of the total charge can be expressed as the sum over variances over all PMTs of the setup for an event:
$\begin{split}\sigma_{Q}^{2}=\sum_{i}^{N_{PMT}}(1+v_{{1}_{i}})\mu_{i}+v_{i}(p)\mu_{i}^{2}=\\ =Q_{0}\frac{1}{N_{PMT}}\sum_{i}^{N_{PMT}}(1+v_{1_{i}})+\\ +Q_{0}^{2}\frac{1}{N_{PMT}}\sum_{i}^{N_{PMT}}v_{i}(p)=\\ =Q_{0}(1+\overline{v_{1}})+Q_{0}^{2}\frac{\overline{v(p)}}{N_{PMT}},\end{split}$ (21)
where $N_{PMT}$ is the total number of (equal) PMTs in the setup. I ignored intrinsic resolution contribution for simplicity and used the following notations for average and relative variance of the involved quantities:
$\overline{v_{1}}\equiv\frac{1}{N_{PMT}}\sum_{i}^{N_{PMT}}v_{1_{i}},$ (22)
$\overline{v(p)}\equiv\frac{1}{N_{PMT}}\sum_{i}^{N_{PMT}}v_{i}(p).$ (23)
Thus, in the case the i-th PMTs exhibit non-uniformity $v(p)_{i}$, in the setup consisting of 4 PMTs the resulting contribution will be a quarter of the average, i.e. $v(p)=\frac{\overline{v(p)}}{4}$, where $\overline{v(p)}=\frac{\sum v(p)_{i}}{4}$.
@@
Figure 5: The fit of the statistical energy resolution data from [6] (see Fig.13 of [6]) with $\frac{1+v_{1}}{LY\cdot f_{\text{NL}}\cdot E}$ function.
At the next stage of the analysis, we observe that the residual energy resolution values for protons, once the statistical component has been subtracted, are at the level of $1\times 10^{-4}$. As the design of the setup is very similar to previous cases, we can expect a similar contribution from the non-uniformity of the response on a single PMT. As shown above, the contribution from the signal summed from four PMTs will correspondingly be reduced by a factor of 4, providing a contribution at the level of $v(p)\simeq 10^{-4}$. Therefore, we can conclude that the contribution of intrinsic resolution for protons is $\simeq 0$ and neglect it in the analysis.
The fit of the proton energy resolution data is presented in Fig. 6. The best-fit value for $v(p)=(1.00\pm 0.25)\times 10^{-4}$, which agrees with the expected 1/4 of the single PMT case.
Figure 6: The fit of the resolution data (see Fig.14 of [6]) for proton beam with constrained statistical term and free $v(p)$. The smearing of the proton beam is taken into account as a penalty term in the $\chi^{2}$.
@@
To test the assumption that the intrinsic resolution contribution for protons is negligible compared to the total resolution, I performed a more sophisticated analysis incorporating both proton and electron datasets.
In this analysis, the same values for light yield ($LY$), $v_{1}$, and $v(p)$ were assumed for both electrons and protons. The beam widths were fixed to the measured values provided in [6]. The $\chi^{2}$-profile for the proton intrinsic resolution ($v_{1_{p}}^{\text{int}}$) was obtained by performing a fit over a grid of $v_{1_{p}}^{\text{int}}$ values for protons, while leaving $v_{1_{e}}^{\text{int}}$ for electrons as a free parameter.
The light yield (LY) was fixed to the value obtained from the statistical data. The $v_{1}$ parameter was constrained to its best-fit value from the same dataset. The $v(p)$ parameter was constrained in the fit to $v(p)=1.00\pm 0.25$, a value derived from a fit to the proton data with both $v(p)$ and $v_{1_{p}}^{\text{int}}$ left free. This value corresponds to a range of $v(p)=3.0\div 5.0$ for a single PMT, which is consistent with values observed in previously considered setups.
No significant bias in the estimate of $v_{1_{e}}^{\text{int}}$ for electrons was observed. Furthermore, the shape of the $\chi^{2}$-profile excludes values of $v_{1_{p}}^{\text{int}}$ for protons above $0.4\times 10^{-4}$ at 90% C.L..
Figure 7: The fit of the data (see Fig.14 of [6]) for electrons from [6] with constrained statistical term and $v(p)$. The smearing of the electron beam is also taken into account as a penalty term in the fit.
@@ 8 Measurements with the LAB+PPO liquid scintillator
Using our model, we calculated the value of $v_{1}^{\text{int}}$: $v_{\text{int}}=\frac{Q_{1}}{Q}v_{1}^{\text{int}}=\frac{LY\cdot 1}{LY\cdot 0.05\cdot f_{\text{NL}}(E,kB)}v_{1}^{\text{int}}\simeq\frac{1}{0.05\cdot 0.9}v_{1}^{\text{int}}$, we get $v_{1}^{\text{int}}=(9.0\pm 2.3)\times 10^{-4}$, which corresponds to an intrinsic resolution of $(3.0\pm 0.4)$% at 1 MeV 444the values in our paper [1] are different: $v_{1}^{\text{int}}=(11\pm 3)\times^{-4}$ and $R_{\text{int}}(\text{1 MeV})=(3.3\pm 0.4)$%, this is the result of a miscalculation, division by 0.9 instead of multiplication.. The value $f_{\text{NL}}(k_{\text{B}},E_{\text{e}}=0.05\text{ MeV})\simeq 0.9$ is obtained from Fig.3 of [4].
Below, we reanalyse the data using the additional measurement details described in the PhD thesis [5].
@@
The fit of the spectrum obtained using a low-intensity light source is shown in Fig.2.29 of [5]. The data in the plot for two-photoelectrons contribution can be used to evaluate it variance as $v_{2}\equiv\frac{p5}{p4}=(\frac{0.0758}{0.4246})^{2}=0.032$, here p4 and p5 are parameters of the fit corresponding to the mean and variance of the double p.e. peak. However, the relative variance of the single photoelectron spectrum is factor 2 larger, i.e. $v_{1}=0.064$, this already contradicts the $v_{1}=0.05$ value used in the analysis.
Other available measurements of the Hamamatsu type R6231-100 PMT show the presence of a strong branch of underamplified signals in its single p.e. spectrum (see Fig.7 from [20]). The values of $v_{1}$, obtained for the two PMT samples in the aforementioned paper, are $v_{1}=0.27$ and $v_{1}=0.28$. This feature is invisible in Fig.2.29 of [5] because of the high threshold set in the measurements. We note that, although the article [20] does not present an estimate of measurement errors, accurately determining the parameter $v_{1}$ is a challenging task for spectra with a significant contribution from underamplified signals due to the parameter’s high sensitivity to spectral shape features.
The presence of a strong underamplified signal branch in the single-electron spectrum will also influence the average SER value: the peak value will shift above 1 p.e., while in the analysis in [5] the position of the single p.e. is chosen practically at the peak value ($q_{\text{peak}}=0.217$ while 1 p.e. corresponds to $\div0.2$1 in $V\cdot ns$ scale [5]). The values of $c$ and $v_{1}$ can be reconstructed using (19) and (20).
@@
The fit was performed using the function (16) with additional term responsible for final energy width of the electrons sample as discussed in section 2:
@@
where $f_{\text{NL}}(E)=f_{\text{NL}}(k_{B}=0.0196,E)$, and $\Delta E$ is the energy smearing in the Ge detector. I assume rectangular distribution of the electron energies withing the energy bin $\Delta E$. The fit returns:
@@
with $\chi^{2}=32.7/26$ and $\Delta E=(2.2\pm 1.2)$ keV, see Fig. 8. The value of $v(p)$ does not depend on the choice of the set of the parameters $v_{1}$ and $c$, so we can conclude that at considered energies $v(p)$ is negligible in this experimental setup, namely it is $v(p)<6.0\times 10^{-4}$ at 90% C.L. as results from the $\chi^{2}$ profile study. The best-fit value of $\Delta E$ is compatible with the stated width of the energy bin on the Ge detector side of 1 keV, the effect of the Ge detector resolution probably contributes to this value as well.
@@
The obtained value $v_{1}^{\text{int}}=(4.8\pm 2.1)\times 10^{-4}$ corresponds to $R_{\text{int}}(\text{1~MeV})=(2.2\pm 0.5)$ %.
@@ 9 The measurement with LAB-based LS using monoenergetic electrons at 1 MeV
In this section, I examine the measurements reported in [8]. In our previous paper [1], I accepted the reported result without critical assessment. However, as the estimated uncertainties seem too low (see Table 3 of the paper [8]), I investigated the reliability of the presented errors. At first glance, it can be seen that the maximum reported uncertainty in the table is 0.110.11%, while the resulting uncertainty is 0.060.06%. This appears to be the result of a miscalculation, as the absolute uncertainties should be summed quadratically.
@@
The PMT used for the measurements was calibrated using the peak value of the charge spectrum, as the reported resolution of a single- photoelectron (about 20%) is unrealistically good. Although there are no plots of SER in the paper, but a very similar measurement was presented by a subgroup in another paper [21], which provides the opportunity to reanalyse the spectrum. The SER spectrum is presented in Fig.9 of the paper [21]. One can observe a cut pedestal and a discrepancy between the mean values of the two fitted Gaussians: $\text{mean}_{1}=123.4$ and $\text{mean}_{2}=264$. If the two Gaussians represent the single electron response and two-electrons responses, then we should have $\text{mean}_{2}=2\cdot\text{mean}_{1}$. It appears that the pedestal position is not being taken into account. A rough estimate can be obtained from $\text{pedestal}=\text{mean}_{1}-(\text{mean}_{2}-\text{mean}_{1})=-17.2$.
@@
In the case of a Poisson distribution of probabilities of detecting a fixed amount of photoelectrons with an average $\mu$, the probability ratio of a two-electron response to a single electron response is given by $\frac{P(2)}{P(1)}=\frac{\mu}{2}$. The ratio of the corresponding areas, calculated from Fig.7, gives us $\frac{P(2)}{P(1)}=0.348$, and hence we can estimate $\mu\simeq 0.695\pm 0.044$. The uncertainty is estimated from the (unfitted) fraction of signals in the SER valley.,
Using the mean and r.m.s. data from the figure, we derive the following: $q_{1}=\frac{(162.5+17.2)\cdot(1-e^{-\mu})}{\mu}=129.3$ ($\overline{q}=162.5$ from Fig.7) and $v(q)=(\frac{96.86}{162.5+17.2})^{2}=0.290$ ($r.m.s.=96.86$ from Fig.7). The calibration parameter $c$ can be calculated using the values from Fig.7: $c=\frac{\text{mean}_{1}}{q_{1}}=0.92$. Similarly, I estimate $v_{1}=0.10\pm 0.01$ using $v(q)=0.290$. The former should be compared to the reported value of $v_{1}=0.04$, which corresponds to a very low 20% SER resolution. Using these values and taking into account the reported LY=3266 p.e., we obtain for statistical contribution of $\delta_{\text{st}}=(1.68\pm 0.08)$% (instead of 1.75%), and for $\delta_{\text{s.p.e.}}=(0.55\pm 0.11)$% (instead of 0.35%). The uncertainties of these values are dominated by the uncertainty in extracting $\mu$ value. Let me point out that the possible contribution of the internal resolution of the LaBr3 detector is neglected in the estimate of systematic uncertainties in [8]; this point requires clarification. In the absence of dedicated measurements, we will take this fact as granted.
The recalculated value of the internal contribution at 0.976 MeV is $R_{\text{int}}(0.976~\text{MeV})=(1.85\pm 0.18)$%. Note the larger uncertainty and the fact that the central value has hardly changed compared to the value reported in [8]. Excluding the double count of the non-uniformity contribution from the final result increases the value of the internal contribution to $R_{\text{int}}(0.976)=(2.06\pm 0.14)$%. The latter value will be used to derive a corrected value of $v_{1}^{\text{int}}$ for this measurement.
@@ 10 Discussion
All data are consistent with our simple model within the measurement precision. Note that in a number of cases, the reported uncertainties were rescaled as they appeared to be overestimated. Furthermore, the data considered do not allow us to draw any conclusions regarding the presence of unaccounted additional fluctuations.
The values of the parameter $v_{1}^{\text{int}}$ obtained in present analysis, along with results derived in [1] from Borexino publications, are compiled in Table 3.
The data indicate an additional broadening of the energy resolution due to intrinsic resolution, contributing approximately $\simeq(1.5\div 2)$% at 1 MeV for all considered scintillators. For the BC408 (EJ200) plastic scintillator, different experiments report IR contributions in the range $\simeq(1.25\div 2.0)$% at 1 MeV. The higher values correspond to measurements around 1 MeV, while lower-energy data yield smaller values. This discrepancy may arise from unaccounted systematics in low-energy measurements, particularly since the results in [6] are free of such systematics. Alternatively, it could suggest different underlying processes in the scintillator response below and above 1 MeV, leading to deviations from the phenomenological model above 1 MeV.
@@
The two Borexino measurements for PC-based LS show good agreement. The value reported in the more recent Borexino publication [11] is likely more precise, as the parameterisation incorporates the energy dependence of the IR contribution. The earlier analysis yields an effective value; converting this into a universal parameter would require an accurate determination of the effective energy corresponding to the tail of the 14C spectrum above the analysis threshold. We further note that this result is unaffected by calibration bias and $v_{1}$ uncertainties, as it was obtained using the PMT occupancy (the number of triggered PMTs per event) rather than the total charge. See [1] for a detailed description of the measurement methodology.
@@
Finally, an experiment with a monoenergetic proton source shows no intrinsic smearing effect, a finding that may be significant for the calibration of organic scintillator detectors.
@@
Table 3: The universal parameter $v_{1}^{\text{int}}$ and the corresponding contribution to the energy resolution at 1 MeV, $R_{\text{int}}(\text{1 MeV}$), for organic scintillators considered in this paper. We add also two measurements from Borexino detector, considered in [1].